# Knowledge-Augmented Multimodal Clinical Rationale Generation for Disease Diagnosis with Small Language Models

### v3

@@ 5 Conclusion and Future Work
We present ClinRaGen, a knowledge-augmented framework that enhances SLMs with LLM-derived reasoning and structured medical knowledge for disease diagnosis and multimodal rationale generation. It introduces a knowledge-augmented attention module that jointly unifies time-series and textual EHRs in the same encoding space while injecting domain knowledge for reliable rationale generation and a sequential multimodal distillation paradigm for transferring LLMs’ reasoning capabilities to SLMs. Extensive evaluations on real world datasets, including quantitative and qualitative analyses, show that ClinRaGen enables SLMs to achieve LLM-comparable performance in disease diagnosis and multimodal rationale generation. This work bridges the performance gap between LLMs and SLMs in clinical tasks. Future research will extend ClinRaGen to a broader range of SLM architectures, datasets, and medical applications.
@@
  • First, although rationale distillation transfers reasoning capabilities from LLMs to SLMs, potential biases in LLM-generated rationales may persist.
@@
Future work will refine knowledge integration, enhance bias mitigation strategies, and extend evaluations to diverse clinical settings.
@@
The datasets utilized in our research are publicly accessible and feature de-identified patient data, accessing these datasets still requires passing the CITI Exam [^4] https://about.citiprogram.org/ and download from Physionet [^5] https://physionet.org/. In addition, this study used the Azure OpenAI service and completed the “opting out of the review process” agreement.
@@ A.1 Lab Test Anomaly Caption
A.2 Prompts for Multimodal Rationale Generation Via ChatGPT
The overall procedure for ChatGPT generating clinical rationale is illustrated in Figure 3. The specific medical note-based rationale prompt and lab test-based rationale prompt are detailed as follows. Medical note-based rationale prompt for ChatGPT,
@@ A.2 Prompts for Multimodal Rationale Generation Via ChatGPT
Lab test-based rationale prompt for ChatGPT, we denote the lab test anomalies as $\bm{T}^{*}$:
@@ A.3 Baseline Details
  • LLaMA: The LLaMA-7B model Touvron et al. (2023), a prominent large language model, employs Reinforcement Learning with Human Feedback (RLHF) and instructional tuning, showcasing its adaptability across diverse NLP tasks. This study applied both zero-shot and fine tuning for disease diagnosis.
@@
  • EHR-KnowGen: As a leading model in EHR multimodal learning, EHR-KnowGen Niu et al. (2024) specializes in generating disease diagnoses. This study excludes external knowledge to maintain a balanced evaluation.
@@ A.5 Rationale Evaluation Metrics
Evaluation scores based on Likert scale: 1). Strongly disagree, 2). Disagree, 3). Neither agree nor disagree, 4). Agree, 5). Strongly agree.
A.6 Rationale Evaluation with BLUE and BERTScore.
Models MIMIC-III MIMIC-IV
BLEU BERTScore BLEU BERTScore
Mistral 0.0163 0.7348 0.0532 0.7625
LLaMA2 0.1441 0.8714 0.2357 0.8808
LLaMA3 0.1641 0.8804 0.2568 0.8919
ClinRaGen 0.2689 0.8972 0.2963 0.9044
Table 5: Rationale evaluation with BLEU and BERTScore.
A.7 Discussion on SLMs Selection
In this section, we discuss our choice of Flan-T5 as the base SLM for our research, focusing on the following aspects: 1). Required CoT ability: Flan-T5 Chung et al. (2024) has been extensively instruction-tuned on numerous datasets and hundreds of tasks, endowing it with strong zero-shot, few-shot, and Chain-of-Thought (CoT) abilities that outperform the original T5 Raffel et al. (2020). In contrast, other SLMs, such as OPT Zhang et al. (2022) and GPT-2 Radford et al. (2019), lack these robust CoT capabilities, which is crucial as a initialization ability for further clinical reasoning distillation. 2). Maximizing SLM potential for practical usage: Although other instruction-finetuned SLMs (e.g., Flan-PaLM) exist, they have substantially larger parameter counts (ranging from 8B to 540B), which is not practical in real world clinical applications and not our target SLMs to investigate. We selected Flan-T5-Small (80M) and Flan-T5-Large (780M) as our base models to maximize the potential of SLMs for accurate disease diagnosis and LLM-comparable multimodal reasoning, while maintaining cost-effectiveness in practical applications. Although we currently use Flan-T5, future work will explore a broader range of SLM architectures to further enhance accuracy.
A.8 Discussion on Teacher LLMs Selection
In this section, we discuss our choice of ChatGPT (GPT-3.5-turbo) as the teacher LLM for our research, focusing on the following aspects: 1). High quality clinical rationales : Although ChatGPT is known for its strong language modeling capabilities, its generated rationales may still contain noise and bias—issues that are critical in precision medicine. To address this, we incorporate external medical domain knowledge and introduce a novel knowledge-augmented attention mechanism during multimodal clinical rationale generation. Our extensive experiments (Section 4.3) show that ClinRaGen effectively mitigates incomplete or incorrect diagnoses and rationales distilled from the teacher LLM, thereby reducing the impact of bias in LLM-generated outputs for SLMs distillation. 2) Test set leakage : The PhysioNet Credentialed Data Use Agreement prohibits the use of MIMIC-series data in public LLMs’ training and applications [^7] https://physionet.org/news/post/gpt-responsible-use, ensuring that test set leakage is not an issue with ChatGPT. Furthermore, the poor performance of ChatGPT under zero-shot prompting (as shown in Table 1) indicates that MIMIC-III and MIMIC-IV data were not used in its training. 3) More powerful LLM for evaluation: While we did not choose the most powerful LLM as our teacher, our current teacher LLM sufficiently enhances SLM capabilities in disease diagnosis and clinical rationale generation. Our evaluations (Sections 4.2 and 4.3), supported by quantitative metrics, comparisons with a superior LLM, and human assessments, confirm the effectiveness of using ChatGPT as the teacher LLM through our multimodal rationale distillation paradigm and knowledge-augmented attention mechanism to improve SLMs’ accuracy in disease diagnosis and modality-consistent rationale generation. In future work, more powerful LLMs can be seamlessly integrated with our method to further enhance evaluation accuracy and robustness.

### v5

@@ 4.3 Rationale Evaluation with LLM-as-a-Judge and Human
Models MIMIC-III MIMIC-IV
BLEU BERTScore BLEU BERTScore
Mistral 0.0163 0.7348 0.0532 0.7625
LLaMA2 0.1441 0.8714 0.2357 0.8808
LLaMA3 0.1641 0.8804 0.2568 0.8919
ClinRaGen 0.2689 0.8972 0.2963 0.9044
Table 4: Rationale evaluation with BLEU and BERTScore.
@@ 5 Conclusion and Future Work
We present ClinRaGen, a knowledge-augmented framework that enhances SLMs with LLM-derived reasoning and structured medical knowledge for disease diagnosis and multimodal rationale generation. It introduces a knowledge-augmented attention module that unifies time-series and textual EHRs within a shared encoding space while incorporating domain knowledge for reliable multimodal rationale generation. Additionally, it presents a sequential multimodal distillation paradigm for transferring the reasoning capabilities of LLMs to SLMs. The extensive evaluations performed on two real-world medical datasets, MIMIC-III and MIMIC-IV, which included both quantitative and qualitative analyses, demonstrate that ClinRaGen allows SLMs to achieve performance comparable to LLMs in disease diagnosis and the generation of multimodal rationales. Furthermore, it can mitigate the bias inherent in distilled rationales from LLMs. This work effectively bridges the performance gap between LLMs and SLMs in clinical tasks. Future research will extend ClinRaGen to a broader range of SLM architectures, LLMs, datasets, and medical applications.
@@
  • First, although knowledge-augmented rationale distillation transfers reasoning capabilities from LLMs to SLMs, some potential biases in LLM-generated rationales may still persist.
@@
Future work will refine knowledge integration, enhance more effective bias mitigation strategies, and extend evaluations to diverse clinical settings.
@@
The datasets utilized in our research are publicly accessible and feature de-identified patient data; accessing these datasets still requires passing the CITI Exam [^4] https://about.citiprogram.org/ and downloading from Physionet [^5] https://physionet.org/. In addition, this study used the Azure OpenAI service and completed the "opting out of the review process" agreement.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No.62432006, No.62276159), Tencent Rhino-Bird Focused Research Program (Value-aligned Credible Large Language Model) and RMGS project (Artificial Intelligence and Big Data Analytics for Social Good).
@@ A.1 Lab Test Anomaly Caption
A.2 Prompts for Multimodal Rationale Generation Via LLMs
The overall procedure for teacher LLMs generating clinical rationale is illustrated in Figure 3. The specific medical note-based rationale prompt and lab test-based rationale prompt are detailed as follows. Medical note-based rationale prompt for LLMs,
@@ A.2 Prompts for Multimodal Rationale Generation Via LLMs
Lab test-based rationale prompt for LLMs, we denote the lab test anomalies as $\bm{T}^{*}$:
@@
Base SLMs Method Distill From ChatGPT Distill From DeepSeek
Micro F1 Macro F1 Micro F1 Macro F1
MIMIC-III
Flan-T5-0.1B Clinical-COT 0.6323 0.6021 0.6338 0.6003
ClinRaGen 0.6410 0.6113 0.6447 0.6121
Flan-T5-0.7B Clinical-COT 0.6398 0.6059 0.6403 0.6072
ClinRaGen 0.6501 0.6196 0.6509 0.6222
OPT-0.1B Clinical-COT 0.4573 0.4099 0.4571 0.4092
ClinRaGen 0.4929 0.4514 0.5009 0.4555
QWEN-2.5-0.5B Clinical-COT 0.5921 0.5400 0.5950 0.5566
ClinRaGen 0.6207 0.5688 0.6219 0.5775
MIMIC-IV
Flan-T5-0.1B Clinical-COT 0.6925 0.6643 0.6933 0.6647
ClinRaGen 0.6989 0.6685 0.6994 0.6712
Flan-T5-0.7B Clinical-COT 0.7055 0.6801 0.7062 0.6803
ClinRaGen 0.7127 0.6893 0.7129 0.6896
OPT-0.1B Clinical-COT 0.5067 0.4547 0.5063 0.4533
ClinRaGen 0.5284 0.4810 0.5315 0.4884
QWEN-2.5-0.5B Clinical-COT 0.6114 0.5882 0.6095 0.5351
ClinRaGen 0.6332 0.6105 0.6406 0.5583
Table 6: Ablation studies on different SLMs and LLMs choice.
@@ A.3 Baseline Details
  • LLaMA: The LLaMA-7B model Touvron et al. (2023), a prominent large language model, employs Reinforcement Learning with Human Feedback (RLHF) and instructional tuning, showcasing its adaptability across diverse NLP tasks. This study applied both zero-shot and fine-tuning for disease diagnosis.
@@
  • EHR-KnowGen: As a leading model in EHR multimodal learning, EHR-KnowGen Niu et al. (2024a) specializes in generating disease diagnoses. This study excludes external knowledge to maintain a balanced evaluation.
@@ A.5 Rationale Evaluation Metrics
Evaluation scores based on Likert scale:
  1. 1.
    Strongly disagree
  2. 2.
    Disagree
  3. 3.
    Neither agree nor disagree
  4. 4.
    Agree
  5. 5.
    Strongly agree
A.6 Discussion on Different SLMs and LLMs
In this section, we discuss the influence of different choices of base SLMs and teacher LLMs for our research, as presented in Table 6. There are several findings: 1). Decoder-Only SLMs Benefit More: Decoder-only models, such as OPTZhang et al. (2022) and QWEN-2.5Qwen et al. (2025), gain an average improvement of 3% over Clinical-COT (single-modal CoT distillation) when using our multimodal CoT distillation method. This boost is attributed to their fully autoregressive nature, which integrates multimodal inputs and outputs within a single processing pipeline, effectively capturing implicit cross-modal dependencies. 2). Encoder-Decoder SLMs Achieve Better Overall Performance: Encoder-decoder models handle longer input sequences more efficiently, since they do not allocate input space to labels, leading to improved disease diagnosis performance, particularly in environments with constrained computing resources. 3). Comparable Performance Across Teacher LLMs: SLMs distilled from DeepSeek and ChatGPT yield similar micro and macro F1 scores in disease diagnosis, underscoring the generality and stability of our approach. This consistency is achieved through carefully designed few-shot instruction templates for LLMs.
Overall, these results indicate that our multimodal CoT distillation mechanism ClinRaGen is a robust and generic method for enhancing diverse SLMs to achieve performance comparable to LLMs in healthcare tasks.